# Effective Community Detection Over Streaming Bipartite Networks (Technical Report)

### v1

@@ 2.4. The Problem of the Community Detection Over Streaming Bipartite Network (CD-SBN)
The CD-SBN Problem: We are now ready to define the problem of detection for communities over a streaming bipartite network upon dynamic updates.
@@
Note that, the CD-SBN problem over streaming bipartite network in Definition 2.8 has two scenarios, one-time snapshot and continuous CD-SBN versions. In particular, the one-time snapshot CD-SBN problem finds CD-SBN communities over a snapshot, $G_{t}$, of a streaming bipartite graph at timestamp $t$, whereas the continuous CD-SBN continuously monitors CD-SBN communities upon edge updates (i.e., $S$) in $G_{t}$, for some registered community constraints in Definition 2.8. The former version can be used to obtain initial CD-SBN community answers when CD-SBN predicates are online specified, and the latter one dynamically maintains CD-SBN community results upon graph updates.
Challenges: The processing of snapshot/continuous CD-SBN queries is quite challenging. One straightforward method to tackle this problem is to enumerate all community candidates and check whether or not these candidates satisfy the CD-SBN constraints. However, this method is quite inefficient. First, the checking of CD-SBN community predicates is very costly and involves the constraints of keywords, graph structural cohesiveness, and user relationship scores. Second, there are many possible CD-SBN community candidates, which incurs high computation costs to enumerate and refine. Third, upon graph updates over a sliding window $W_{t}$, it is also challenging to incrementally maintain the dynamically changing graph $G_{t}$ and efficiently update the CD-SBN query answers. Table 1 shows the commonly used symbols and their descriptions in this paper.
3. The CD-SBN Processing Framework
Figure 1 illustrates a CD-SBN processing framework for answering snapshot/continuous CD-SBN queries, which consists of three components, initialization, graph incremental maintenance, and CD-SBN query processing.
To support the CD-SBN stream processing, the initialization component prepares some auxiliary data over initial bipartite graph $G_{0}$ at timestamp 00, including a keyword bit vector $v.BV$ for hashing keywords in item vertices, wedge-score-related data $X_{u_{i},u_{l}}$ and $Y_{u_{i},u_{l}}$ for the butterfly score calculation, support upper bound $ub\_sup(e_{u,v})$, and synopses, $Syn$ (lines 1-4).
At each timestamp $t$, the sliding window $W_{t}$ includes a new item $p_{t}$ (i.e., the edge insertion) and removes an expired item $p_{t-s}$ (i.e., the edge deletion). Therefore, the graph incremental maintenance component will compute an incremental update of wedge-score-related data, $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$ (given by Eqs. (5.2) and  (5.2), respectively), and obtain the updated wedge-score-related data $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ at timestamp $t$ (lines 5-6). Similarly, for those affected edges $e^{\prime}$ (due to edge insertion/deletion), we will also update support upper bounds $ub\_sup(e^{\prime})$ (line 7) and synopsis $Syn$ (line 8).
For the CD-SBN query processing component, the snapshot CD-SBN algorithm traverses the synopsis $Syn$ to retrieve candidate communities $g$, by applying our proposed pruning strategies, and refines the resulting candidate communities $g$ to obtain/return actual CD-SBN answers in $R$ (lines 9-12). Moreover, the continuous CD-SBN algorithm incrementally maintains CD-SBN community answers in $R$ by removing those communities (with the expired edge) and inserting new CD-SBN communities (due to the edge insertion) (lines 13-15).
@@ 3. The CD-SBN Processing Framework
3 calculate the wedge-score-related data $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ (as given in Eq. (4.5)) for each pair of user vertices $u_{i},u_{j}\in U(G_{0})$
4 compute the support upper bound $ub\_sup(e_{u,v})$ for each edge $e_{u,v}\in E(G_{0})$
@@
8       compute $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$ (as given in Eqs. (5.2) and  (5.2), resp.) and obtain the updated $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$
@@
13             traverse the synopsis $Syn$ by applying keyword, support, and score upper bound pruning strategies to retrieve candidate communities $g$
@@ 4. Pruning Strategies
This section details the community-level pruning strategies (see Section 3) to reduce the search space of the CD-SBN problem.
@@ 4.1. Keyword Pruning
Based on the CD-SBN problem definition (as given Definition 2.8), any item vertex $v_{i}$ in the lower layer of a candidate bipartite subgraph $g$ should contain at least one query keyword in $Q$. In the sequel, we present an effective keyword pruning, which filters out item vertices $v_{i}$ in subgraph $g$ that do not contain any query keywords in $Q$.
@@ 4.2. Support Pruning
Discussions on How to Obtain the Upper Bound of Support $\mathbf{ub\_sup(e_{u,v})}$. In the offline phase, the upper bound of support needs to be computed to enable the support pruning. Since part of item vertices in data graph $G_{t}$ will be filtered out by the keyword pruning, where the rest of the vertices induce a subgraph $g$, the number of butterflies containing the edge $e_{u,v}$ in $g$ is smaller than or equal to that in $G_{t}$. Thus, the support of $e_{u,v}$ in $G_{t}$ can be considered as a supper upper bound $ub\_sup(e_{u,v})$ in subgraph $g$ defined below.
(4) $\vspace{-1.5ex}ub\_sup(e_{u,v})=\sum_{\forall v_{x}\in N(u)-\{v\}}{|N(v)\cap N(v_{x})-\{u\}|},\vspace{-0.5ex}$
@@ 4.3. Layer Size Pruning
Since the support of a bipartite subgraph is closely related to the subgraph size, we can utilize filter out a candidate community $g$ with small sizes of upper/lower layers (i.e., $|U(g)|$ and $|L(g)|$), with respect to the support threshold $k$.
@@ 4.5. Score Upper Bound Pruning
Proof.
Given a candidate user vertex $u_{i}$ and a user vertex $u_{j}\in U(g)$, since $ub\_score_{u_{i},u_{j}}(g)$ is a score upper bound, it holds that $score_{u_{i},u_{j}}(g)\leq ub\_score_{u_{i},u_{j}}(g)$. From the assumption of the lemma, if $ub\_score_{u_{i},u_{j}}(g)<\sigma$ holds, by the inequality transition, we have $score_{u_{i},u_{j}}(g)\leq ub\_score_{u_{i},u_{j}}(g)<\sigma$, violating the score constraint of the $(k,r,\sigma)$-bitruss (given in Definition 2.7). Thus, the candidate user vertex $u_{i}$ can be safely pruned, which completes the proof. ∎
@@
$=$ $\frac{1}{2}\left(\left(\sum_{\forall v\in L(G_{t})}{\min\{w_{u_{i},v},w_{u_{j},v}\}}\right)^{2}-\sum_{\forall v\in L(G_{t})}\left(\min\{w_{u_{i},v},w_{u_{j},v}\}\right)^{2}\right)$
@@ 5. Auxiliary Data Initialization and Incremental Maintenance
This section details the initialization stage (lines 1-4 of Algorithm 1) that computes auxiliary data over an initial bipartite graph $G_{0}$ in Section 5.1, and constructs a synopsis, $Syn$, for such auxiliary data in Section 5.3, as well as incremental maintenance of auxiliary data and synopsis (lines 5-8 of Algorithm 1) in Sections 5.2 and  5.4, respectively.
@@ 5.1. Auxiliary Data Initialization
To facilitate efficient online community detection, we will compute some auxiliary data over the initial graph $G_{0}$ in the initialization phase, which can enable our proposed pruning strategies (as mentioned in Section 4).
@@
  • (User Keyword Bit Vector) a keyword bit vector, $u.BV$, of each user vertex $u$, which is a bit-OR of bit vectors $v.BV$ for all $u$’s neighbor item vertices $v\in N(u)$ (i.e., $u.BV=\bigvee_{\forall v\in N(u))}v.BV$);
@@
  • (Aggregated Wedge Scores) the summed wedge score, $X_{u_{i},u_{j}}$ $=\sum_{\forall v\in L(G_{t})}$ $w_{\angle(u_{i},v,u_{j})}$, and the summation over the squares of wedge scores, $Y_{u_{i},u_{j}}=$ $\sum_{\forall v\in L(G_{t})}(w_{\angle(u_{i},v,u_{j})})^{2}$, for each user pair $(u_{i},u_{j})\in U(G_{0})\times U(G_{0})$ and item vertices $v\in N(u_{i})\cap N(u_{j})$.
@@ 5.2. Incremental Maintenance of Auxiliary Data
In this subsection, we illustrate how to incrementally maintain auxiliary data of graph $G_{t}$ for each user vertex, upon changes of the update items in the sliding window $W_{t}$. As mentioned in Section 2.1, the sliding window $W_{t}$ at a new timestamp $t$ (from $W_{t-1}$) needs to insert a new item $p_{t}$ and remove an expired old item $p_{t-s}$. Upon such updates, we need to incrementally compute the edge weights in the graph $G_{t}$, as well as the auxiliary data in the synopsis $Syn$.
Maintenance of Keyword Bit Vectors: When an edge $e_{u_{i},v_{a}}$ is inserted or deleted in the bipartite graph $G_{t}$, we need to update the keyword bit vector $u_{i}.BV$. Specifically, for the insertion of a new edge $e_{u_{i},v_{a}}$, the new keyword bit vector is given by $u_{i}.BV\bigvee v_{a}.BV$; for the deletion of an expired edge $e_{u_{i},v_{a}}$, we recompute the keyword bit vector $u_{i}.BV=\bigvee_{\forall v\in N(u_{i})}v.BV$.
Maintenance of Support Upper Bounds: At timestamp $t$, due to the insertion of $p_{t}$ or the deletion of $p_{t-s}$, an edge $e_{u_{i},v_{a}}$ can be added or deleted, respectively. Consequently, we need to update the support upper bounds of the edges in all butterflies $\Join(u_{i},u_{j},v_{a},v_{b})$ containing the edge $e_{u_{i},v_{a}}$. When inserting a new edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we increase the support upper bound $ub\_sup(e)$ by 1. Similarly, when deleting an edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we decrease the support upper bound $ub\_sup(e)$ by 1.
Maintenance of User Relationship Scores: Since the user relationship score $score_{u_{i},u_{j}}(G_{t})$ is calculated using $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ (as given by Eq. (4.5)), we maintain these two terms incrementally. Specifically, we compute the differences, $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$, of $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$, respectively, between timestamps $(t-1)$ and $t$. Upon updating the edge $e_{u_{i},v_{a}}$ at timestamp $t$, the edge weight $w_{u_{i},v_{a}}$ has an increase/decrease of 1. Accordingly, the wedge weight $w_{\angle(u_{i},v_{a},u_{j})}$ ($=min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$) is updated with $w_{\angle(u_{i},v_{a},u_{j})}$ $+$ $\lambda_{\angle(u_{i},v_{a},u_{j})}$, where $\lambda_{\angle(u_{i},v_{a},u_{j})}$$\in$$\{-1,0,1\}$ (depending on the change of $min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$).
@@
$\Delta X_{u_{i},u_{j}}$
$=$ $X_{u_{i},u_{j}}^{(t)}-X_{u_{i},u_{j}}^{(t-1)}$
$=$ $\left(\sum_{\forall v\in L(G),v\neq v_{a}}{w_{\angle(u_{i},v,u_{j})}}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})\right)-\sum_{\forall v\in L(G)}{w_{\angle(u_{i},v,u_{j})}}$
$=$ $\lambda_{\angle(u_{i},v_{a},u_{j})},$
where $X_{u_{i},u_{j}}^{(t)}$ and $X_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $X_{u_{i},u_{j}}$ at timestamps $t$ and $(t-1)$, respectively.
From Eq. (5.2), for any user pair $(u_{i},u_{j})$, the increase/decrease of $X_{u_{i},u_{j}}$ is equal to that of $w_{\angle(u_{i},v_{a},u_{j})}$ (i.e., $\lambda_{\angle(u_{i},v_{a},u_{j})}$).
$\Delta Y_{u_{i},u_{j}}$
$=$ $Y_{u_{i},u_{j}}^{(t)}-Y_{u_{i},u_{j}}^{(t-1)}$
$=$ $\left(\sum_{\forall v\in L(G),v\neq v_{a}}\left(w_{\angle(u_{i},v,u_{j})}\right)^{2}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})^{2}\right)$
@@
$=$ $2\lambda_{\angle(u_{i},v_{a},u_{j})}w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})}^{2},$
where $Y_{u_{i},u_{j}}^{(t)}$ and $Y_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $Y_{u_{i},u_{j}}$ at timestamps $t$ and $(t-1)$, respectively.
As given in Eq. (5.2), the increment/decrement of $Y_{u_{i},u_{j}}$ can be computed via $w_{\angle(u_{i},v_{a},u_{j})}$ and $\lambda_{\angle(u_{i},v_{a},u_{j})}$.
Time Complexity Analysis: When a new edge $e_{u_{i},v_{a}}$ is inserted, the updated keyword bit vector is given by $u_{i}.BV\bigvee v_{a}.BV$, which takes $O(1)$ cost; when an edge $e_{u_{i},v_{a}}$ is deleted, we can re-compute the keyword bit vector $u_{i}.BV=\bigvee_{\forall v\in N(u_{i})}v.BV$, which has the time complexity $O(|N(u_{i})|)$.
For an updated edge $e_{u_{i},v_{a}}$, we update the support upper bound of edges connected to $v_{a}$ and $v_{b}$, where $v_{b}\in N(u_{i})\cap N(u_{j})-\{v_{a}\},u_{j}\in N(v_{a})$. Thus, the worst-case time cost is $O(|N(v_{a})|\cdot|N(u_{i})|)$.
Upon the update of edge $e_{u_{i},v_{a}}$, for any user pair $(u_{i},u_{j})$, the time complexity of updating $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ is given by $O(1)$. The worst-case time complexity is given by $O(|N(v_{a})|)$, where $N(v_{a})$ is a set of user neighbors $u_{j}$ of the item vertex $v_{a}$.
Therefore, the total time complexity of the graph increment maintenance is given by $O(|N(v_{a})|\cdot|N(u_{i})|)$.

### v2

@@ 2.4. The Problem of the Community Detection Over Streaming Bipartite Network (CD-SBN)
The CD-SBN Problem: We are now ready to define the problem of community detection over streaming bipartite network.
@@
Note that the CD-SBN problem has two scenarios: one-time snapshot and continuous CD-SBN. In particular, the one-time snapshot CD-SBN problem finds CD-SBN communities over a snapshot, $G_{t}$, of a streaming bipartite graph at timestamp $t$, whereas the continuous CD-SBN continuously monitors CD-SBN communities upon edge updates (i.e., $S$) in $G_{t}$, for some registered community constraints in Definition 2.8. The former version can be used to obtain initial CD-SBN community answers when CD-SBN predicates are online-specified, and the latter dynamically maintains CD-SBN community results upon graph updates.
Challenges: The processing of snapshot and continuous CD-SBN queries is challenging. One straightforward method to tackle the snapshot problem is to enumerate all community candidates and check whether they satisfy the CD-SBN constraints. However, this method is quite inefficient because of the costly checking of CD-SBN community predicates, including how to efficiently verify the constraints of keywords, graph structural cohesiveness, and user relationship scores. For the continuous CD-SBN problem, a straightforward method is to answer the snapshot query at each timestamp $t$ upon graph updates over a sliding window $W_{t}$. Nonetheless, there are many possible CD-SBN community candidates, which incur high computation costs to enumerate and refine. Therefore, it is also challenging to incrementally maintain the dynamically changing graph $G_{t}$ and efficiently update the CD-SBN query answers.
Table 1 lists the commonly used symbols and their descriptions in this paper.
@@
3. The CD-SBN Processing Framework
Figure 1 illustrates a framework for answering snapshot/continuous CD-SBN queries, which consists of three components, initialization, graph incremental maintenance, and CD-SBN query processing.
To support the CD-SBN stream processing, the initialization component prepares some auxiliary data over the initial bipartite graph $G_{0}$ at timestamp 0, including a keyword bit vector $v.BV$ for hashing keywords in item vertices, wedge-score-related data for the butterfly score calculation, support upper bound $ub\_sup(e_{u,v})$, and a synopsis $Syn$ (lines 1-4).
At each timestamp $t$, the sliding window $W_{t}$ includes a new item $p_{t}$ (i.e., the edge insertion) and removes an expired item $p_{t-s}$ (i.e., the edge deletion). Therefore, a graph incremental maintenance component will compute an incremental update of wedge-score-related data and obtain the updated wedge-score-related data at timestamp $t$ (lines 5-6). Similarly, for those affected edges $e^{\prime}$ (due to edge insertion/deletion), we will update support upper bounds $ub\_sup(e^{\prime})$ (line 7) and synopsis $Syn$ (line 8).
For the CD-SBN query processing component, the snapshot CD-SBN algorithm traverses the synopsis $Syn$ to retrieve candidate communities $g$ by applying our proposed pruning strategies, and refines the resulting candidate communities $g$ to obtain actual CD-SBN answers in $R$ (lines 8-11). Moreover, the continuous CD-SBN algorithm incrementally maintains CD-SBN community answers in $R$ by removing communities with expired edges and inserting new CD-SBN communities due to edge insertion (lines 12-14).
@@ 3. The CD-SBN Processing Framework
3 calculate the wedge-score-related data for each pair of user vertices $u_{i},u_{j}\in U(G_{0})$ compute the support upper bound $ub\_sup(e_{u,v})$ for each edge $e_{u,v}\in E(G_{0})$
@@
7  compute the increment of wedge-score-related data and update the data for each pair of user vertices $u_{i},u_{j}\in U(G_{t})$
@@
12     traverse the synopsis SynSyn by applying keyword, support, and score upper bound pruning strategies to retrieve candidate communities $g$
@@ 4. Pruning Strategies
This section outlines the community-level pruning strategies (see Section 3) for reducing the search space of the CD-SBN problem.
@@ 4.1. Keyword Pruning
Based on the CD-SBN problem definition (Definition 2.8), any item vertex $v_{i}$ in the lower layer of a candidate bipartite subgraph $g$ should contain at least one query keyword in $Q$. In the sequel, we present an effective keyword pruning, which filters out item vertices $v_{i}$ in subgraph $g$ that do not contain any query keywords in $Q$.
@@ 4.2. Support Pruning
Discussions on How to Obtain the Upper Bound of Support $\mathbf{ub\_sup(e_{u,v})}$. In the offline phase, the upper bound on support must be computed to enable support pruning. Since part of the item vertices in the data graph $G_{t}$ will be filtered out by the keyword pruning, where the rest of the vertices induce a subgraph $g$, the number of butterflies containing the edge $e_{u,v}$ in $g$ is smaller than or equal to that in $G_{t}$. Thus, the support of $e_{u,v}$ in $G_{t}$ can be considered as an upper bound $ub\_sup(e_{u,v})$ in the subgraph $g$ defined below.
(4) $ub\_sup(e_{u,v})=\sum_{\forall v_{x}\in N(u)-\{v\}}{|N(v)\cap N(v_{x})-\{u\}|},$
@@ 4.3. Layer Size Pruning
Since the support of a bipartite subgraph is closely related to the subgraph size, we can filter out a candidate community $g$ with small sizes of upper/lower layers (i.e., $|U(g)|$ and $|L(g)|$), with respect to the support threshold $k$.
@@ 4.5. Score Upper Bound Pruning
$=$ $\frac{1}{2}\left(\left(\sum_{\forall v\in L(G_{t})}\hskip-8.61108pt{\min\{w_{u_{i},v},w_{u_{j},v}\}}\right)^{2}\hskip-4.30554pt-\hskip-8.61108pt\sum_{\forall v\in L(G_{t})}\hskip-8.61108pt\left(\min\{w_{u_{i},v},w_{u_{j},v}\}\right)^{2}\right)$
@@ 5. Auxiliary Data Initialization and Incremental Maintenance
This section details the initialization stage (lines 1-4 of Algorithm 1) that computes auxiliary data over an initial bipartite graph $G_{0}$ in Section 5.1, constructs a synopsis, $Syn$, for such auxiliary data in Section 5.3, and incremental maintenance of auxiliary data and synopsis (lines 5-8 of Algorithm 1) in Sections 5.2 and  5.4, respectively.
@@ 5.1. Auxiliary Data Initialization
To facilitate efficient online community detection, we will compute auxiliary data over the initial graph $G_{0}$ during the initialization phase, which will enable our proposed pruning strategies (as mentioned in Section 4).
@@
  • (Aggregated Wedge Scores) the summed wedge score, $X_{u_{i},u_{j}}$ $\sum_{\forall v\in L(G_{t})}$ $w_{\angle(u_{i},v,u_{j})}$, and the summation over the squares of wedge scores, $Y_{u_{i},u_{j}}=$ $\sum_{\forall v\in L(G_{t})}(w_{\angle(u_{i},v,u_{j})})^{2}$, for each user pair $(u_{i},u_{j})\in U(G_{0})\times U(G_{0})$ and item vertices $v\in N(u_{i})\cap N(u_{j})$.
@@ 5.2. Incremental Maintenance of Auxiliary Data
In this subsection, we illustrate how to incrementally maintain auxiliary data for each user vertex in graph $G_{t}$ upon changes to the update items in the sliding window $W_{t}$. As mentioned in Section 2.1, the sliding window $W_{t}$ at a new timestamp $t$ (from $W_{t-1}$) needs to insert a new item $p_{t}$ and remove an expired old item $p_{t-s}$. Upon such updates, we need to incrementally compute the edge weights in the graph $G_{t}$, as well as the auxiliary data in the synopsis $Syn$.
Maintenance of Support Upper Bounds: At timestamp $t$, due to the insertion of $p_{t}$ or the deletion of $p_{t-s}$, an edge $e_{u_{i},v_{a}}$ can be added or deleted, respectively. Consequently, we need to update the support upper bounds of the edges in all butterflies $\Join(u_{i},u_{j},v_{a},v_{b})$ containing the edge $e_{u_{i},v_{a}}$. When inserting a new edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we increase the support upper bound ub_sup(e)ub\_sup(e) by 1. Similarly, when deleting an edge $e_{u_{i},v_{a}}$, for each edge $e$ in $\Join(u_{i},u_{j},v_{a},v_{b})$, we decrease the support upper bound ub_sup(e)ub\_sup(e) by 1.
Maintenance of User Relationship Scores: Since the user relationship score $score_{u_{i},u_{j}}(G_{t})$ is calculated using $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ (as given by Eq. (4.5)), we maintain these two terms incrementally. Specifically, we compute the differences, $\Delta X_{u_{i},u_{j}}$ and $\Delta Y_{u_{i},u_{j}}$, of $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$, respectively, between timestamps (t1)(t-1) and $t$. Upon updating the edge $e_{u_{i},v_{a}}$ at timestamp $t$, the edge weight $w_{u_{i},v_{a}}$ has an increase/decrease of 1. Accordingly, the wedge weight $w_{\angle(u_{i},v_{a},u_{j})}$ ($min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$) is updated with $w_{\angle(u_{i},v_{a},u_{j})}$ ++ $\lambda_{\angle(u_{i},v_{a},u_{j})}$, where $\lambda_{\angle(u_{i},v_{a},u_{j})}\in$$\{-1,0,1\}$ (depending on the change of $min\{w_{u_{i},v_{a}},w_{u_{j},v_{a}}\}$).
@@
$\hskip-8.61108pt\Delta X_{u_{i},u_{j}}$
$\hskip-8.61108pt=$ $\hskip-8.61108ptX_{u_{i},u_{j}}^{(t)}-X_{u_{i},u_{j}}^{(t-1)}$
$\hskip-8.61108pt=$ $\hskip-8.61108pt\left(\sum_{\forall v\in L(G),v\neq v_{a}}\hskip-17.22217pt{w_{\angle(u_{i},v,u_{j})}}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})\right)-\hskip-8.61108pt\sum_{\forall v\in L(G)}\hskip-8.61108pt{w_{\angle(u_{i},v,u_{j})}}\hskip-17.22217pt$
$\hskip-8.61108pt=$ $\hskip-8.61108pt\lambda_{\angle(u_{i},v_{a},u_{j})},$
where $X_{u_{i},u_{j}}^{(t)}$ and $X_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $X_{u_{i},u_{j}}$ at timestamps $t$ and (t1)(t-1), respectively.
$\hskip-8.61108pt\Delta Y_{u_{i},u_{j}}$
$\hskip-8.61108pt=$ $\hskip-8.61108ptY_{u_{i},u_{j}}^{(t)}-Y_{u_{i},u_{j}}^{(t-1)}$
$\hskip-8.61108pt=$ $\hskip-8.61108pt\left(\sum_{\forall v\in L(G),v\neq v_{a}}\left(w_{\angle(u_{i},v,u_{j})}\right)^{2}+(w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})})^{2}\right)$
@@
$\hskip-8.61108pt=$ $\hskip-8.61108pt2\lambda_{\angle(u_{i},v_{a},u_{j})}w_{\angle(u_{i},v_{a},u_{j})}+\lambda_{\angle(u_{i},v_{a},u_{j})}^{2},$
where $Y_{u_{i},u_{j}}^{(t)}$ and $Y_{u_{i},u_{j}}^{(t-1)}$ are the wedge-score-related data $Y_{u_{i},u_{j}}$ at timestamps $t$ and (t1)(t-1), respectively.
As given in Eq. (5.2) and Eq. (5.2), for any user pair $(u_{i},u_{j})$, the increase/decrease of $X_{u_{i},u_{j}}$ is equal to $\lambda_{\angle(u_{i},v_{a},u_{j})}$, and the increment/decrement of $Y_{u_{i},u_{j}}$ can be computed via $w_{\angle(u_{i},v_{a},u_{j})}$ and $\lambda_{\angle(u_{i},v_{a},u_{j})}$.
Time Complexity Analysis: For an updated edge $e_{u_{i},v_{a}}$, we update the support upper bound of edges connected to $v_{a}$ and $v_{b}$, where $v_{b}\in N(u_{i})\cap N(u_{j})-\{v_{a}\},u_{j}\in N(v_{a})$. Thus, the worst-case time cost is $O(|N(v_{a})|\cdot|N(u_{i})|)$. Upon the update of edge $e_{u_{i},v_{a}}$, for any user pair $(u_{i},u_{j})$, the time complexity of updating $X_{u_{i},u_{j}}$ and $Y_{u_{i},u_{j}}$ is given by $O(1)$. The worst-case time complexity is given by $O(|N(v_{a})|)$, where $N(v_{a})$ is a set of user neighbors $u_{j}$ of the item vertex $v_{a}$. Therefore, the total time complexity of the graph increment maintenance is given by $O(|N(v_{a})|\cdot|N(u_{i})|)$.